# CMM-Math: A Chinese Multimodal Math Dataset To Evaluate and Enhance the Mathematics Reasoning of Large Multimodal Models

### v2

@@ Introduction
CogVLM2 19B 25.85 21.94 25.22 24.76 28.85 24.68 26.54 29.79 44.44 19.05 33.79 23.36 45.45 25.58
InternLM-VL 11.1B 19.82 21.94 26.38 20.68 21.15 12.99 16.67 19.15 0.00 9.52 20.48 12.15 27.27 13.95
CogVLM2 (3-Shot) 19B 31.21 30.97 28.12 32.57 32.69 24.68 30.25 29.79 22.22 25.40 36.18 30.84 45.45 27.91
InternLM-VL (3-Shot) 11.1B 25.09 29.03 30.14 24.76 15.38 27.27 25.93 12.77 0.00 26.98 24.57 17.76 9.09 25.58
@@
Qwen-VL - 49.91 49.68 48.12 60.59 44.23 50.65 54.94 21.28 11.11 50.79 42.66 36.45 63.64 51.16
Gemini - 41.88 36.77 37.97 55.37 26.92 37.66 48.15 31.91 33.33 26.98 37.54 32.71 54.55 25.58
GPT-4o - 29.02 26.13 28.41 35.67 26.92 20.78 32.10 25.53 33.33 28.57 24.57 21.96 18.18 37.21
Qwen-VL (3-Shot) - 64.91 70.00 63.19 74.43 50.00 66.23 69.75 21.28 44.44 73.02 54.61 57.01 63.64 53.49
Gemini (3-Shot) - 41.65 38.71 35.07 54.07 19.23 31.17 51.85 27.66 44.44 44.44 36.18 30.84 54.55 44.19
GPT-4o (3-Shot) - 65.98 62.9 59.71 82.57 46.15 66.23 72.84 40.43 77.78 68.25 59.04 47.66 81.82 55.81
Math-LMM 8.26B 32.10 26.13 29.86 36.64 28.85 32.47 35.19 27.66 33.33 30.16 35.15 28.04 36.36 25.58
Mean accuracy of LMMs - 38.86 37.65 37.47 45.65 30.94 35.89 42.20 26.11 31.31 36.65 36.80 30.80 45.45 35.09
Table 5: Comparison of model performances in accuracy across various mathematical subjects. Subjects: Alg: algebra, AnaG: analytic geometry, Ari: arithmetic, CombG: combinatorial geometry, Comb: combinatorics, Cnt: counting, Desc: deseriptive geometry, GrphT: graph theory, Log: logic, MetG: metric geometry, SolG: solid geometry, Stat: statistics, TransG: transformation geometry. The first and second highest accuracy of LMMs are marked in red and blue, respectively.
Models Size Overall LV1 LV2 LV3 LV4 LV5 LV6 LV7 LV8 LV9 LV10 LV11 LV12
Open-source LMMs
CogVLM2 19B 25.85 38.00 36.36 21.15 29.38 20.39 24.32 27.32 29.33 24.9 23.02 25.37 22.73
InternLM-VL 11.1B 19.82 18.00 15.91 20.19 12.99 6.58 28.83 23.90 26.0 26.97 13.06 22.69 20.63
CogVLM2 (3-Shot) 19B 31.21 37.00 34.09 40.38 29.94 28.95 35.14 34.63 32.67 28.22 28.52 32.84 25.52
InternLM-VL (3-Shot) 11.1B 25.09 22.00 32.95 27.88 16.95 11.84 27.93 25.37 26.67 24.9 24.05 26.87 31.82
Closed-source LMMs
Qwen-VL - 49.91 70.00 62.50 56.73 54.80 46.71 67.57 46.83 43.33 41.08 45.7 48.96 46.85
Gemini - 41.88 65.00 65.91 58.65 51.41 46.05 56.76 39.02 39.33 34.85 35.40 35.52 29.72
GPT-4o - 29.02 46.00 43.18 38.46 33.90 32.89 32.43 27.32 24.67 28.63 19.59 27.76 23.78
Qwen-VL (3-Shot) - 64.91 60.00 76.14 72.12 69.49 57.24 71.17 64.39 58.67 52.70 66.67 68.06 67.83
Gemini (3-Shot) - 41.65 69.00 64.77 52.88 41.24 46.05 55.86 38.54 41.33 39.83 33.33 33.73 34.97
GPT-4o (3-Shot) - 65.98 80.00 84.09 75.00 81.92 69.74 72.07 68.78 61.33 56.85 52.23 66.57 59.44
Math-LMM 8.26B 32.10 32.00 44.32 37.50 35.59 32.24 35.14 32.20 28.67 32.37 30.58 31.94 26.22
Mean accuracy of LMMs - 38.86 48.82 50.93 45.54 41.60 36.24 46.11 38.94 37.45 35.57 33.83 38.21 35.41
Table 6: Comparison of model performances in accuracy across different levels. The levels from 1 to 12 correspond to grades from primary school to high school.
Models Size Overall Alg AnaG Ari CombG Comb Cnt Desc GrphT Log MetG SolG Stat TransG
Open-source LMMs
CogVLM2 19B 2.82 2.6 2.44 3.21 2.81 2.44 2.29 2.05 2.41 2.89 2.91 2.41 2.71 3.36
InternLM-VL 11.1B 4.48 4.51 4.02 5.24 2.69 4.28 4.08 2.11 3.82 4.75 4.16 3.78 5.14 3.07
CogVLM2 (3-Shot) 19B 2.72 2.53 2.62 2.94 2.42 2.59 2.65 1.90 2.12 2.84 2.68 2.51 2.57 3.14
InternLM-VL (3-Shot) 11.1B 4.35 4.58 4.14 5.08 2.91 4.25 3.73 2.06 3.06 4.44 3.91 3.65 4.71 2.93
Closed-source LMMs
Qwen-VL - 6.50 6.66 5.80 7.41 4.30 6.07 5.84 4.65 3.82 6.16 6.12 5.86 6.29 5.07
Gemini - 6.02 5.65 5.29 6.72 5.72 5.15 5.32 5.17 5.47 5.30 6.06 5.71 6.00 5.05
GPT-4o - 7.94 8.15 7.28 8.65 6.23 7.89 7.70 6.22 7.41 7.56 7.65 7.20 8.00 7.14
Qwen-VL (3-Shot) - 6.21 6.52 6.14 6.73 4.74 5.44 5.45 5.74 5.59 6.06 5.96 5.81 4.86 4.81
Gemini (3-Shot) - 5.89 5.72 5.41 6.45 5.62 5.49 5.36 5.34 6.18 5.33 5.78 5.41 6.14 5.02
GPT-4o (3-Shot) - 7.85 7.99 7.07 8.61 6.12 7.74 7.51 6.31 8.24 7.52 7.46 7.22 8.57 6.90
Math-LMM 8.26B 2.46 2.19 2.23 2.76 2.40 1.90 1.84 2.41 2.06 2.62 2.40 2.53 1.43 2.43
Mean score of LMMs - 5.20 5.19 4.77 5.80 4.18 4.84 4.71 4.00 4.56 5.04 5.01 4.74 5.13 4.45
Table 7: Comparison of model performances in GPT-4o score across various mathematical subjects.
Models Size Overall LV1 LV2 LV3 LV4 LV5 LV6 LV7 LV8 LV9 LV10 LV11 LV12
Open-source LMMs
CogVLM2 19B 2.82 2.96 3.66 3.52 2.99 2.29 2.9 2.46 3.05 2.37 2.43 2.45 2.43
InternLM-VL 11.1B 4.48 4.05 5.54 5.27 4.9 4.2 5.12 4.0 4.03 2.6 4.13 4.41 4.35
CogVLM2 (3-Shot) 19B 2.72 2.73 3.53 3.18 2.78 2.3 2.66 2.35 2.97 2.31 2.46 2.48 2.74
InternLM-VL (3-Shot) 11.1B 4.35 3.95 5.46 5.06 4.62 3.77 4.78 3.96 4.05 3.08 4.11 4.31 4.36
Closed-source LMMs
Qwen-VL - 6.50 5.91 7.30 7.54 7.11 6.61 7.02 6.1 5.91 4.73 6.01 6.18 6.17
Gemini - 6.02 6.30 6.85 6.83 6.46 5.75 6.40 5.67 6.46 5.01 5.42 4.96 5.62
GPT-4o - 7.94 7.44 8.70 8.76 8.34 8.04 8.30 7.83 7.65 6.44 7.67 7.48 7.74
Qwen-VL (3-Shot) - 6.21 5.66 6.70 6.43 6.18 6.12 6.91 6.00 5.72 5.50 5.67 6.57 6.51
Gemini (3-Shot) - 5.89 5.83 6.54 6.47 6.06 5.44 6.26 5.51 6.59 5.36 5.77 5.21 5.44
GPT-4o (3-Shot) - 7.85 7.27 8.53 8.63 8.45 8.09 8.36 7.87 7.65 6.06 7.48 7.22 7.46
Math-LMM 8.26B 2.46 2.27 3.17 2.95 2.97 2.4 1.83 2.16 2.67 2.16 2.13 2.26 2.11
Mean score of LMMs - 5.20 4.94 6.00 5.88 5.53 5.00 5.50 4.90 5.16 4.15 4.84 4.87 4.99
Table 8: Comparison of model performances in GPT-4o score across different levels. The levels from 1 to 12 correspond to primary to high school grades. The maximum score is 10.
Model Input Overall FQA GPS MWP TQA VQA ALG ARI GEO LOG NUM SCI STA
Heuristics baselines
Random chance - 17.9 18.2 21.6 3.8 19.6 26.3 21.7 14.7 20.1 13.5 8.3 17.2 16.3
Frequent guess - 26.3 22.7 34.1 20.4 31.0 24.6 33.1 18.7 31.4 24.3 19.4 32.0 20.9
Large Multimodal Models (LMMs)
IDEFICS-9B-Instruct $Q$, $I$ 19.8 21.6 21.1 6.5 25.9 24.0 22.1 15.0 19.8 18.9 9.9 24.6 18.1
mPLUG-Owl-LLaMA-7B $Q$, $I$ 22.2 22.7 23.6 10.2 27.2 27.9 23.6 19.2 23.9 13.5 12.7 26.3 21.4
miniGPT4-LLaMA-2-7B $Q$, $I$ 23.1 18.6 26.0 13.4 30.4 30.2 28.1 21.0 24.7 16.2 16.7 25.4 17.9
LLaMA-Adapter-V2-7B $Q$, $I$ 23.9 21.2 25.5 11.3 32.3 31.8 26.3 20.4 24.3 24.3 13.9 29.5 18.3
LLaVAR $Q$, $I$ 25.2 21.9 25.0 16.7 34.8 30.7 24.2 22.1 23.0 13.5 15.3 42.6 21.9
InstructBLIP-Vicuna-7B $Q$, $I$ 25.3 23.1 20.7 18.3 32.3 35.2 21.8 27.1 20.7 18.9 20.4 33.0 23.1
LLaVA-LLaMA-2-13B $Q$, $I$ 26.1 26.8 29.3 16.1 32.3 26.3 27.3 20.1 28.8 24.3 18.3 37.3 25.1
Multimodal Bard $Q$, $I$ 34.8 26.0 47.1 29.6 48.7 26.8 46.5 28.6 47.8 13.5 14.9 47.5 33.0
GPT-4V (Playground) $Q$, $I$ 49.9 43.1 50.5 57.5 65.2 38.0 53.0 49.0 51.0 21.6 20.1 63.1 55.8
Math-LMM (Ours) $Q$, $I$ 34.9 25.3 46.6 46.2 34.2 24.6 43.1 30.3 45.6 16.2 27.1 28.6 25.9
Human
Human performance $Q$, $I$ 60.3 59.7 48.4 73.0 63.2 55.9 50.9 59.2 51.4 40.7 53.8 64.9 63.9
Table 9: Accuracy scores on the testmini subset of MathVista. Input: $Q$: question, $I$: image. FQA: figure question answering, GPS: geometry problem solving, MWP: math word problem, TQA: textbook question answering, VQA: visual question answering. Mathematical reasoning types: ALG: algebraic reasoning, ARI: arithmetic reasoning, GEO: geometry reasoning, LOG: logical reasoning, NUM: numeric commonsense, SCI: scientific reasoning, STA: statistical reasoning.
Random Chance 7.17 1.5 11.9 7.1 9.7 4.8 6.0 22.1 1.1 7.6 0.6 9.4 6.7 8.2 8.6 13.0 7.1
Closed-source LMMs (Text + Image)
Qwen-VL-Plus 10.72 11.3 17.9 14.3 12.7 4.8 10.5 15.4 8.9 14.3 11.6 6.4 10.0 14.3 6.9 8.7 11.31
Qwen-VL-Max 15.59 10.7 19.1 20.0 16.9 12.5 17.9 16.4 12.2 21.0 13.3 14.2 19.8 11.5 20.7 13.0 17.3
Gemini Pro 17.66 15.1 10.7 20.7 20.1 11.9 7.5 20.2 21.1 16.8 19.1 19.0 20.0 14.3 13.8 17.4 20.8
GPT-4V 22.76 27.3 32.1 35.7 21.1 16.7 13.4 22.1 14.4 16.8 22.0 22.2 20.9 23.8 24.1 21.7 25.6
Open-source LMMs (Text + Image)
LLaVA-v1.5-7B 8.52 7.0 7.1 10.7 7.1 4.8 10.5 7.7 10.0 9.2 15.6 10.2 9.8 5.3 8.6 4.4 4.8
SPHINX (V2) 9.70 6.7 7.1 12.9 7.5 7.7 6.0 9.6 16.7 10.1 11.0 11.8 12.5 8.2 8.6 8.7 6.0
ShareGPT4V-7B 10.53 5.5 3.6 12.9 10.1 4.8 7.5 11.5 14.4 10.9 16.2 11.8 12.3 9.8 15.5 17.4 11.3
LLaVA-v1.5-13B 11.12 7.0 14.3 14.3 9.1 6.6 6.0 13.5 5.6 13.5 10.4 12.6 14.7 11.5 13.8 13.0 10.7
ShareGPT4V-13B 11.88 7.5 15.5 16.4 10.7 8.9 9.0 11.5 8.9 7.6 11.6 13.0 17.4 10.3 8.6 8.7 12.5
SPHINX-MoE 14.18 7.8 17.9 14.3 15.6 9.5 11.9 12.5 15.6 12.6 16.2 15.6 17.8 13.5 12.1 8.7 16.1
InternLM-XComposer2-VL 14.54 9.3 15.5 12.1 15.3 11.3 10.5 14.4 22.2 19.3 19.7 15.6 15.0 11.9 15.5 26.1 15.5
Math-LMM (Ours) 11.58 7.3 8.3 10.7 14.0 7.1 7.4 16.4 12.2 9.2 14.5 10.6 14.9 9.0 8.6 26.1 16.7
@@
Human (testmini) 75.66 57.9 79.0 100.0 100.0 47.4 94.7 89.5 63.2 63.2 36.8 52.6 73.7 89.5 89.5 100.0 73.7
Table 10: Comparison of model performances on MATH-V across various mathematical subjects. Subjects: Alg: algebra, AnaG: analytic geometry, Ari: arithmetic, CombG: combinatorial geometry, Comb: combinatorics, Cnt: counting, DescG: descriptive geometry, GrphT: graph theory, Log: logic, Angle: metric geometry - angle, Area: metric geometry - area, Len: metric geometry-length, SolG: solid geometry, Stat: statistics, Topo: topology, TransG: transformation geometry.
Implementation Details.
We employ a lightweight DFN5B-H-14+ (Fang et al. 2023) as the vision encoder. This model is trained on a large amount of high-quality paired image-text data using Contrastive Language Image Pre-training (CLIP) and demonstrates superior performance at a comparable scale of parameters. The image encoding module contains only 0.6 billion parameters, which enables our model to perform inference deployment on a single 4090 GPU. Table 4 presents the parameter details of the Math-LMM model.
Evaluating SOTA LMMs on CMM-Math
We evaluate the performance of several state-of-the-art large multimodal models (LMMs) over CMM-MATH in terms of accuracy (Table 5 and Table 6) and GPT-4o score (Table 7 and Table 8). We also report the performance of various subjects and levels to analyze the LMMs’ abilities of multimodal mathematical reasoning.
From the results, we observe the following findings. First, our model, Math-LMM, surpasses other open-source LMMs in overall accuracy, achieving a score of 32.10 with only 8.26 billion parameters, compared to 31.21 for other models. Second, LMMs excel in arithmetic and statistics but show limited proficiency in geometry. We can observe that the mean accuracy of all LMMs across all subjects is 38.86, with LMMs achieving an accuracy exceeding 45 in arithmetic and statistics subjects. In contrast, in geometry, including analytical geometry, combinatorial geometry, descriptive geometry, metric geometry, and transformation geometry, the accuracy of LMMs is below 38.86, with the worst performance in descriptive geometry, where the accuracy rate is 26.11. Third, LMMs perform well on primary school-level problems, but their performance declines at the high school level. As indicated in the last row of Table 6 and Table 8, the mean accuracy of LMMs shows an overall downward trend. Specifically, for instance, when employing three-shot prompting with GPT-4o, the model achieves an accuracy of 84.08 on level 2 problems but only 52.23 on level 10 problems. This phenomenon can also be more intuitively observed in Figure 3. Fourth, GPT-4o and Qwen-VL demonstrate the best performance on the CMM-MATH in most cases. GPT-4o performs optimally on simple and moderate difficulty levels, whereas Qwen-VL excels on the most challenging levels, including levels 10, 11, and 12. Fifth, it is interesting that most LLMs are good at choice and yes-no problems while the performance on fill-in-the-blank and analysis problems is limited. The possible reason is that most training datasets are choice and yes-no problems. The LMMs are not good at giving the solution step by step. Furthermore, we noticed that Math-LMM performed poorly, which may be due to the relatively less training corpus and the smaller model scale, leading to weaker language expression and analytical reasoning capabilities, resulting in lower scores.
Evaluating Math-LMM
@@
To evaluate the effectiveness of our Math-LMM, we conduct experiments on MATHVISTA and MATH-V (Table 9 and Table 10). These two datasets are widely used to evaluate the capability of large multimodal models in solving mathematical problems.
MATHVISTA
@@
We selected LMMs used in the MathVista (Lu et al. 2024) experiments for comparison with Math-LMM, including IDEFICS-9B-Instruct (Laurençon et al. 2023), mPLUG-Owl-LLaMA-7B (Ye et al. 2024), miniGPT4-LLaMA-2-7B (Zhu et al. 2023), LLaMA-Adapter-V2-7B (Gao et al. 2023b), LLaVAR (Zhang et al. 2024b), InstructBLIP-Vicuna-7B (Dai et al. 2023), LLaVA-LLaMA-2-13B (Liu et al. 2023a), Multimodal Bard (Google 2023), and GPT-4V (Playground) (OpenAI 2023). Table 9 reveals that our model, Math-LMM, achieved the second-best overall score of 34.9 on MathVista (Lu et al. 2024), ranking just below GPT-4V. Furthermore, Math-LMM achieved the highest score of 27.1 in the numeric commonsense task (NUM) and the second-highest score of 46.2 in the math word problem task (MWP). These outcomes suggest that Math-LMM demonstrates commendable capabilities in addressing English multimodal mathematical questions. Detailed comparative analyses of other models can be found in MathVista (Lu et al. 2024).
MATH-V
@@
We also conducted experiments using the LMMs employed in MATH-V (Wang et al. 2024), including Qwen-VL-Plus (Bai et al. 2023b), Qwen-VL-Max (Bai et al. 2023b), Gemini Pro (Team et al. 2024), GPT-4V (OpenAI 2023), LLaVa-v1.5-7B (Liu et al. 2023a), SPHINX (Lin et al. 2023), ShareGPT-4V-7B (Chen et al. 2023), LLaVa-v1.5-13B (Liu et al. 2023a), ShareGPT-4V-13B (Chen et al. 2023), SPHINX-MoE (Lin et al. 2023), and InternLM-XComposer2-VL (Dong et al. 2024).
In Table 10, compared to open-source models of the same parameter scale, LLaVa-v1.5-7B (Liu et al. 2023a), SPHINX (Lin et al. 2023), and ShareGPT-4V-7B (Chen et al. 2023), Math-LMM achieved the best performance, scoring 11.58. Against larger-parameter-scale open-source models, Math-LMM outperformed LLaVa-v1.5-13B (Liu et al. 2023a) but underperformed compared to ShareGPT-4V-13B (Chen et al. 2023), SPHINX-MoE (Lin et al. 2023), and InternLM-XComposer2-VL (Dong et al. 2024). Additionally, Math-LMM achieved the best performance among all models on the topology subject. These results indicate that Math-LMM also has a certain level of competitiveness on English multimodal mathematical problems.
Conclusion

### v3

@@ 5.2. Evaluating SOTA LMMs on CMM-Math
LLaVA-v1.5-7B 8.52 7.00 7.10 10.70 7.10 4.80 10.50 7.70 10.00 9.20 15.60 10.20 9.80 5.30 8.60 4.40 4.80
SPHINX (V2) 9.70 6.70 7.10 12.90 7.50 7.70 6.00 9.60 16.70 10.10 11.00 11.80 12.50 8.20 8.60 8.70 6.00
ShareGPT4V-7B 10.53 5.50 3.60 12.90 10.10 4.80 7.50 11.50 14.40 10.90 16.20 11.80 12.30 9.80 15.50 17.40 11.30
LLaVA-v1.5-13B 11.12 7.00 14.30 14.30 9.10 6.60 6.00 13.50 5.60 13.50 10.40 12.60 14.70 11.50 13.80 13.00 10.70
ShareGPT4V-13B 11.88 7.50 15.50 16.40 10.70 8.90 9.00 11.50 8.90 7.60 11.60 13.00 17.40 10.30 8.60 8.70 12.50
SPHINX-MoE 14.18 7.80 17.90 14.30 15.60 9.50 11.90 12.50 15.6 12.60 16.20 15.60 17.80 13.50 12.10 8.70 16.10
InternLM-VL 14.54 9.30 15.50 12.10 15.30 11.30 10.50 14.40 22.20 19.30 19.70 15.6 15.00 11.90 15.50 26.10 15.50
Math-LMM (Ours 7B) 11.58 7.30 8.30 10.70 14.00 7.10 7.40 16.40 12.20 9.20 14.50 10.60 14.90 9.00 8.60 26.10 16.70
Math-LMM (Ours 72B) 17.53 10.70 28.60 15.00 20.10 11.30 11.90 15.40 16.70 21.00 22.50 18.40 20.00 15.60 20.70 8.70 19.60
@@
Qwen-VL-Plus 10.72 11.30 17.90 14.30 12.70 4.80 10.50 15.40 8.90 14.30 11.60 6.40 10.00 14.30 6.90 8.70 11.31
Qwen-VL-Max 15.59 10.70 19.10 20.00 16.90 12.50 17.90 16.40 12.20 21.00 13.30 14.20 19.80 11.50 20.70 13.00 17.30
Gemini Pro 17.66 15.10 10.70 20.70 20.10 11.90 7.50 20.20 21.10 16.80 19.10 19.00 20.00 14.30 13.80 17.40 20.80
GPT-4V 22.76 27.30 32.10 35.70 21.10 16.70 13.40 22.10 14.40 16.80 22.00 22.20 20.90 23.80 24.10 21.70 25.60
@@
Human (testmini) 75.66 57.90 79.00 100.00 100.00 47.40 94.70 89.50 63.20 63.20 36.80 52.60 73.70 89.50 89.50 100.00 73.70
Disparity Between Accuracy and GPT-4o Score. When comparing the performance of LMMs in both accuracy and GPT-4o score, we observe that closed-source models maintain consistent performance while open-source models show inconsistencies. Specifically, GPT-4o consistently ranks highest, while Gemini ranks lowest. CogVLM2 has higher accuracy than InternLM-VL but is lower on the GPT-4o score. This discrepancy may arise from differences in the training data distribution, where the datasets might focus more on choice and yes-no questions while lacking materials for problem-solving and analytical tasks. Moreover, this inconsistency could also result from insufficient training data, leading to models that obtain correct answers but struggle to provide high-quality analytical processes. This highlights that our CMM-Math, unlike other math test sets that offer only choice questions, provides a more comprehensive evaluation of models’ analytical and problem-solving abilities.
Zero-shot versus 3-shot Scenarios. All few-shot models perform better than zero-shot prompting in terms of accuracy. This indicates that few-shot prompting helps models obtain correct answers. Conversely, regarding the GPT-4o score, most models perform better with zero-shot prompting than with few-shot prompting, except for the LLaVA-v1.5 and LLaVA-v1.6-mistral models. This suggests that few-shot prompting does not necessarily lead to higher quality outputs and analyses. Our view is that the content of few-shot prompts may influence how models conduct their own analysis and reasoning, potentially leading to poorer performance.
Attention to SOTA LMMs
We separately provide an analysis of the performance of closed-source and open-source LMMs.
Closed-source LMMs. We observe that all closed-source models significantly outperform open-source models. Both in terms of accuracy and GPT-4o score, GPT-4o and Qwen-VL demonstrate the best performance on the CMM-Math in most cases. However, we note that GPT-4o performs optimally on simple and moderately difficult levels, while Qwen-VL excels on the most challenging levels, including levels 10, 11, and 12. This suggests that Qwen-VL may have a stronger advantage in handling more complex problems.
Open-source LMMs. In terms of accuracy, our Math-LMM (72B) and Qwen2-VL-Instruct (3-Shot) achieve the best and second-best performances, with scores of 48.57 and 46.29, respectively. On the most challenging levels 10, 11, and 12, Math-LMM (72B) achieves the best results, indicating that Math-LMM has a stronger advantage in handling more complex problems. Furthermore, regarding the GPT-4o score, we notice that Math-LMM achieved only ordinary results, while Qwen2-VL-Instruct obtains the best result of 5.26. This could be because the training data for Math-LMM is still limited, leading to weaker language expression and analytical reasoning capabilities.
5.3. Evaluating Math-LMM
@@
To evaluate the effectiveness of our Math-LMM, we conduct experiments on MATHVISTA (Lu et al., 2024) and MATH-V (Wang et al., 2024b) (Table 8 and Table 9). These two datasets are widely used to evaluate the capability of large multimodal models in solving English mathematical problems.
Main Results over MATHVISTA
@@
We select LMMs used in the MathVista (Lu et al., 2024) experiments for comparison with Math-LMM, including IDEFICS-9B-Instruct (Laurençon et al., 2023), mPLUG-Owl-LLaMA-7B (Ye et al., 2024), miniGPT4-LLaMA-2-7B (Zhu et al., 2023), LLaMA-Adapter-V2-7B (Gao et al., 2023a), LLaVAR (Zhang et al., 2024b), InstructBLIP-Vicuna-7B (Dai et al., 2023), LLaVA-LLaMA-2-13B (Liu et al., 2024b), Multimodal Bard (Google, 2023), and GPT-4V (Playground) (OpenAI, 2023).
In the results of open-source models shown in Table 8, our Math-LMM achieves the best (36.3) and suboptimal (34.9) results with 72B and 7B versions. Across five problem tasks, Math-LMM (7B) obtains the best performance on two tasks and Math-LMM (72B) achieves the best performance on three tasks and the second-best on one task. Furthermore, our Math-LMM almost achieves the best performance across all types of mathematical reasoning, except for scientific reasoning, where it achieved a suboptimal performance. Additionally, in numeric commonsense problems, both parameter versions of our Math-LMM perform exceptionally well, even surpassing closed-source models such as Multimodal Bard (Google, 2023) and GPT-4V (Playground). These outcomes suggest that Math-LMM demonstrates commendable capabilities in addressing English multimodal mathematical questions. Detailed comparative analyses of other models can be found in MathVista (Lu et al., 2024).
Main Results over MATH-V
@@
We also compare our Math-LMM with existing LLMs over MATH-V (Wang et al., 2024b), including Qwen-VL-Plus (Bai et al., 2023b), Qwen-VL-Max (Bai et al., 2023b), Gemini Pro (Team et al., 2024), GPT-4V (OpenAI, 2023), LLaVa-v1.5-7B (Liu et al., 2024b), SPHINX (Lin et al., 2023), ShareGPT-4V-7B (Chen et al., 2023), LLaVa-v1.5-13B (Liu et al., 2024b), ShareGPT-4V-13B (Chen et al., 2023), InternLM-XComposer2-VL (Dong et al., 2024), and SPHINX-MoE (Lin et al., 2023).
In Table 9, our Math-LMM (72B) achieves the best performance among in open-source models. Although the Math-LMM (7B) does not achieve the second-best performance, compared to other open-source models of the same parameter scale, such as LLaVa-v1.5-7B (Liu et al., 2024b), SPHINX (Lin et al., 2023), and ShareGPT-4V-7B (Chen et al., 2023), Math-LMM (7B) still achieves the best performance. This may be because MATH-V (Wang et al., 2024b) is a more challenging test set than MATHVISTA (Lu et al., 2024), where model parameter scale has a greater impact on performance. Meanwhile, among open-source models, Math-LMM almost achieves the best performance across all 16 subjects, except for the subjects of arithmetic and graph theory. Notably, compared to closed-source models, Math-LMM also achieves the best performance among all models on the subjects of logic, metric geometry-angle, and topology. These results indicate that Math-LMM also has a certain level of competitiveness on harder English multimodal mathematical problems.
6. Conclusion